3. Apple farmers are always concerned with the weight of the apples they produce. A farmer samples 26 apples from her trees. The sample resulted in a mean weight of 86 grams and standard deviation of 6.2 grams. Obtain a 90% CI for the standard deviation of the weights of all apples this farmer produces..
Answer #1

Here s = 6.2 and n = 26
df = 26 - 1 = 25

α = 1 - 0.9 = 0.1
The critical values for α = 0.1 and df = 25 are ꭓ^2(1-α/2,n-1) = 14.6114 and ꭓ^2(α/2,n-1) = 37.6525

(n − 1)sos (n - 1)s V xázs xá/2 xias

CI = (sqrt(25*6.2^2/37.6525) , sqrt(25*6.2^2/14.6114))
CI = (5.0520 , 8.1099)